Florida Development, Inc.'s (FDI) free cash flow during the year just ended was $75 million, and FCF is expected to grow at constant rate of 6.50% in the future. If the weighted average cost of capital is 13% what is the FDI's value of operations, in millions?
Answer #1

Value of operations=FCF1/(WACC-Growth Rate)

=(75*1.065)/(0.13-0.065)

=(79.875/0.065)

which is equal to

=$1228.85 million(Approx).